(*) Find the cigenvalues and eigenfunctions for the following problem: - = A, 0 < x <l, y(0) = 2y(1), y(0) =yl), where I >
Answer #1

oence - y - dy zo Gloja 2 910) y o ylle y = ARAB = B = 2 Al +2B BE2A Y 10) = 24W y62A = y) - 0 तर valu43zi elsen vectar macove to d = ar so y+ a²y zo mata you) = so a cosax te Sinax y (0) = 24 (1) c = 2 (acosal to sinal) C. (1-2 cosal) - czasinalf its corresponding Eigen function is 1903 - es cas ana x + ce soin 23a